OpenStudy (anonymous):

Toshe is sorting beads by color. She has 6 different colors of beads. There are 42 beads of each color. How many beads does Toshe have in all? A. 7 B. 48 C. 242 D. 252 will fan and medal!!!

OpenStudy (anonymous):

She has how many beads of different colors?

OpenStudy (anonymous):

mmmmm

OpenStudy (anonymous):

let me think

OpenStudy (anonymous):

6 different beads with 42 beads of each colour 42+42+42+42+42+42 = 6*42

OpenStudy (anonymous):

6*42 = 252 right?

OpenStudy (anonymous):

i got 80% and Mastered !!